How would you solve a problem like this.
Is this ordered pair a solution of the system of linear equations.
2x+y=3 (0,3)
x-2y=-1
Please show work.

Found 2 solutions by richard1234, Cintchr:
Answer by richard1234(7193)

Ordered pair is a solution if and only if it satisfies the given equation.

Therefore, to see if (0,3) is a solution to 2x+y = 3, replace x = 0, y = 3. If you obtain a true statement, it is a solution.

Repeat with the other equation. If (0,3) satisfies both equations, then it is a solution to the system.

Answer by Cintchr(481)

Plug in 0 for x and 3 for y in BOTH equations.
.
.
2(0)+ 3 = 3
0 + 3 = 3
3 = 3 TRUE
.
.
0 - 2(3) = -1
-6 = -1
-6 = -1 FALSE
.
.
(0,3) is NOT a solution to this system on linear equations.